Convert 50.0 rev to radians





1
Expert's answer
2022-01-10T09:15:37-0500

Explanations & Calculations


  • Radians and Degrees are units of measurements in circular dimensions.
  • There are "\small 360" degrees for a full round and that in Radians is "\small 2\pi".

"\qquad\qquad \small \text{Round\,=}\,360^0=2\pi^c"

  • Therefore, for "\small 50" revolutions there are "\small 50\times360^0=18000^0" degrees and "\small 50\times2\pi=100\pi=314.16" radians.